OpenStudy (anonymous):

Please can someone help me with this using pseudocode and algorithm to compute the model: A = √(s-b)(s-b)(s-c) S = a+b+c/2 with explanation

OpenStudy (anonymous):

step 1: Define a function (I am using this in the most general sense) ---lets call it CalculateArea ---which accepts inputs of type integer s, b and c.

OpenStudy (anonymous):

The BODMAS/BIDMAS/BIRDMAS rule tells you: do what is in the brackets first, so Step 2. subtract the value b from s and store in a temporary variable called A1 subtract the value b from s and store in a temporary variable called A2 subtract the value c from s and store in a temporary variable called A3 You will then need to multiply these values according to your formula so: Step 3: multiply A1, A2 and A3 and store the result in a temporoary variable called M1 Finally, you will need to get the square root of M1, the result of multiplying these together Step 4: get the square root of M1 and assign this to the variable A Finally you will need to print out the value Step5: Print the value of A to the screen. leave out my comments and just include the steps, and you will have the pseudocode for the first equation.

OpenStudy (anonymous):

Thanks for the answer........what about the algorithm of the equation?

OpenStudy (anonymous):

From the pseudocode, we can deduce the algorithm as follows, (remember an algorithm is just a series of steps). Imagine you are trying to tell a computer how to do this. Algorithm: create a function called CalculateArea which accepts 3 integers called s,b,c A1 = s-b A2 = s-b A3 = s-c M1 = A1 x A2 x A3 A= sqrt (M1) Print A;

OpenStudy (mathmate):

@ubah Isn't the Hero's formula meant to be: \(\color{red}{s} = \color{red}{(a}+b+c\color{red}{)}/2\) \(A = \color{red}{\sqrt{\color{black}{(s-\color{red}{a})(s-b)(s-c)}}}\) In programming, - upper/lower case count - brackets count - names of variables count - order of statements counts The compiler is not very forgiving, it does exactly what you tell it to do.